When the melons were weighed in pairs each pair weighed more than 100 pounds.
smith's and jone's melons together weighed 110 pounds
jone's and Squash's melons together weighed 104 pounds
smith's and squash's melons together weighed 122 pounds
whose melon was the heaviest and how much did it weigh?

Answer by checkley71(8403)

SM+J=100
J+SQ=104 SUBTRACTING EQ #2 FROM EQ#1 WE GET
---------
SM-SQ=-4 ADDING THIS TO EQ #3 WE GET
SM+SQ=122
----------
2SM=118 OR SM=59 THEN SQ=122-59 OR SQ=63 & J+63=104 OR J=41 & 59+41=100 100=100

SM=59, SQ=63 & J=41.